OpenStudy (katherinesmith):

What is the equation of a parabola with a vertex at (1, -1) and a directrix y = 4.

OpenStudy (dumbcow):

 \[y = a(x-1)^{2} -1\] \[a = -\frac{1}{4p}\]

OpenStudy (katherinesmith):

thats not any of my choices

OpenStudy (dumbcow):

haha well do you understand what i did? what value did you plug in for "p" ?